# ProtocoLLM: Automatic Evaluation Framework of LLMs on Domain-Specific Scientific Protocol Formulation Tasks

### v1

@@ Abstract
Automated generation of scientific protocols executable by robots can significantly accelerate scientific research processes. Large Language Models (LLMs) excel at Scientific Protocol Formulation Tasks (SPFT), but the evaluation of their capabilities rely on human evaluation. Here, we propose a flexible, automatic framework to evaluate LLMs’ capability on SPFT: ProtocoLLM [^1] The dataset and code are available here.. This framework prompts the target model and GPT-4 to extract pseudocode from biology protocols using only predefined lab actions and evaluates the output of target model using Llam-Eval, the pseudocode generated by GPT-4 serving as a baseline and Llama-3 acting as the evaluator. Our adaptable prompt-based evaluation method, Llam-Eval, offers significant flexibility in terms of evaluation model, material, criteria, and is free of cost. We evaluate GPT variations, Llama, Mixtral, Gemma, Cohere, and Gemini. Overall, we find that GPT and Cohere is a powerful scientific protocol formulators. We also introduce Bioprot 2.0, a dataset with biology protocols and corresponding pseudocodes, which can aid LLMs in formulation and evaluation of SPFT. Our work is extensible to assess LLMs on SPFT across various domains and other fields that require protocol generation for specific goals.
@@ 1 Introduction
Figure 1: Overview of the ProtocoLLM Framework. A protocol containing a title, descriptions, step-by-step instructions, and predefined biology lab actions is given to both a target model and GPT-4 for pseudocode generation. Then, Llama-3 evaluates these outputs considering the target model’s pseudocode as the prediction ($\hat{y}$) and GPT-4’s as a baseline ($y$).
@@
Here, we propose an evaluation framework that evaluates the capabilities of LLMs in SPFT: ProtocoLLM (Figure 1).
@@
    We propose ProtocoLLM: a flexible, automatic framework for evaluating LLMs on SPFT using domain knowledge and Llam-Eval.
@@ 3 Methods
The ProtocoLLM framework can evaluate the capability of LLMs on SPFT in three steps (Figure 2): (1) prompt the target LLM to generate pseudocode based on the given protocol, (2) repeat previous step for GPT-4, and (3) Llam-Eval for evaluation. To utilize this framework, we curated protocols in biology (Section 3.1), predefined actions performed in biology labs (Section 3.2), prompted LLMs for pseudocode generation (Section 3.3), and prompted Llama-3 for evaluation (Llam-Eval) (Section 3.6).
Figure 2: The ProtocoLLM Framework.
@@ 3.4 Metrics and Evaluation
This approach is not constrained by the output structure of the target models, eliminates the need for manual annotation efforts during the parsing process as required in reference-based metrics, enables a comprehensive evaluation, and thereby makes ProtocoLLM significantly more flexible and automatic.
@@ 5 Conclusion
We introduce ProtocoLLM, a flexible and automatic framework designed to evaluate LLMs’ capabilities on Scientific Protocol Formulation Tasks (SPFT). This framework prompts the target model and GPT-4 to extract pseudocode from biology protocols using only predefined lab actions, then evaluates the target model’s output using Llam-Eval, with the GPT-4 generated pseudocode as a baseline and Llama-3 as the evaluator. Our prompt-based evaluation method, Llam-Eval, provides significant flexibility in terms of evaluation models, materials, criteria, and is free of cost. We assess various models, including GPT variants, Llama, Mixtral, Gemma, Cohere, and Gemini, and find GPT and Cohere to be particularly effective in formulating scientific protocols. Additionally, we present Bioprot 2.0, a dataset containing biology protocols and corresponding pseudocodes, which supports LLMs in the formulation and evaluation of SPFT. Our work is extensible to the assessment of LLMs on SPFT across various domains and other fields that require protocol generation for specific goals.
@@ 6 Limitations
We recognize several limitations. The predefined actions may not encompass all actions performed in a biology labs. The definitions of predefined actions may be incomplete. To precisely define an action, it is necessary to define not only the function but also the function arguments. The number of protocols in Bioprot 2.0 may be insufficient for evaluation purposes. The performance of ProtocoLLM may decline outside of biology. Addressing this requires redefining domain-specific actions and exploring other LLMs for diverse fields. Future work should investigate these cross-disciplinary implications. LLMs are continuously evolving due to regular updates. The LLMs used for evaluation in this work might become unavailable in the future. Upgraded versions of LLMs may result in performance degradation and metrics may differ from those obtained using previous models. Due to selecting Llama-3 as the evaluator, our results may be susceptible to its biases and hallucinations. The outcomes when evaluated with models other than Llama-3 are unknown. Future work should investigate the outcomes using different LLMs as an evaluator. Using an API of LLMs as an evaluator, such GPT, is often not free of charge and can be costly. We used GPT-4 generated responses as a baseline; however, it may not accurately represent the ground truth. Future work should explore the implications of employing alternative resources (e.g., manually annotated pseudocodes, responses generated by other models) as the baseline. We observed basic actions classified as NoAction in minor cases. It has been reported that GPT prefers outputs from LLMs, which also produced our evaluation materials including all ground truth and target pseudocodes. This can potentially influence the scores. The four criteria mentioned in G-Eval may not sufficiently fulfill the role of evaluating protocols where real-world validation is crucial. Also, applying these criteria originally designed for summarization tasks may be inappropriate for evaluating SPFT. Even if the protocol pseudocode is successfully synthesized, real-world experiments may fail depending on the person performing the protocol or the condition of the physical equipment, especially in cases that are more complex than stem cell culture or require delicate manual work and experience.
@@ Acknowledgements
We would like to thank Karim Md.Adnan for assisting us with the action defining process.

### v2

# ProtoMed-LLM: An Automatic Evaluation Framework for Large Language Models in Medical Protocol Formulation

@@ Abstract
Automated generation of scientific protocols executable by robots can significantly accelerate scientific research processes. Large Language Models (LLMs) excel at Scientific Protocol Formulation Tasks (SPFT), but the evaluation of their capabilities rely on human evaluation. Here, we propose a flexible, automatic framework to evaluate LLMs’ capability on SPFT: ProtoMed-LLM [^1] The dataset and code are available here.
*Corresponding Authors
. This framework prompts the target model and GPT-4 to extract pseudocode from biology protocols using only predefined lab actions and evaluates the output of target model using Llam-Eval, the pseudocode generated by GPT-4 serving as a baseline and Llama-3 acting as the evaluator. Our adaptable prompt-based evaluation method, Llam-Eval, offers significant flexibility in terms of evaluation model, material, criteria, and is free of cost. We evaluate GPT variations, Llama, Mixtral, Gemma, Cohere, and Gemini. Overall, we find that GPT and Cohere is a powerful scientific protocol formulators. We also introduce Bioprot 2.0, a dataset with biology protocols and corresponding pseudocodes, which can aid LLMs in formulation and evaluation of SPFT. Our work is extensible to assess LLMs on SPFT across various domains and other fields that require protocol generation for specific goals.
@@ 1 Introduction
Figure 1: Overview of the ProtoMed-LLM Framework. A protocol containing a title, descriptions, step-by-step instructions, and predefined biology lab actions is given to both a target model and GPT-4 for pseudocode generation. Then, Llama-3 evaluates these outputs considering the target model’s pseudocode as the prediction ($\hat{y}$) and GPT-4’s as a baseline ($y$).
@@
Here, we propose an evaluation framework that evaluates the capabilities of LLMs in SPFT: ProtoMed-LLM (Figure 1).
@@
    We propose ProtoMed-LLM: a flexible, automatic framework for evaluating LLMs on SPFT using domain knowledge and Llam-Eval.
@@ 3 Methods
The ProtoMed-LLM framework can evaluate the capability of LLMs on SPFT in three steps (Figure 2): (1) prompt the target LLM to generate pseudocode based on the given protocol, (2) repeat previous step for GPT-4, and (3) Llam-Eval for evaluation. To utilize this framework, we curated protocols in biology (Section 3.1), predefined actions performed in biology labs (Section 3.2), prompted LLMs for pseudocode generation (Section 3.3), and prompted Llama-3 for evaluation (Llam-Eval) (Section 3.6).
Figure 2: The ProtoMed-LLM Framework.
@@ 3.4 Metrics and Evaluation
This approach is not constrained by the output structure of the target models, eliminates the need for manual annotation efforts during the parsing process as required in reference-based metrics, enables a comprehensive evaluation, and thereby makes ProtoMed-LLM significantly more flexible and automatic.
@@ 5 Conclusion
We introduce ProtoMed-LLM, a flexible and automatic framework designed to evaluate LLMs’ capabilities on Scientific Protocol Formulation Tasks (SPFT). This framework prompts the target model and GPT-4 to extract pseudocode from biology protocols using only predefined lab actions, then evaluates the target model’s output using Llam-Eval, with the GPT-4 generated pseudocode as a baseline and Llama-3 as the evaluator. Our prompt-based evaluation method, Llam-Eval, provides significant flexibility in terms of evaluation models, materials, criteria, and is free of cost. We assess various models, including GPT variants, Llama, Mixtral, Gemma, Cohere, and Gemini, and find GPT and Cohere to be particularly effective in formulating scientific protocols. Additionally, we present Bioprot 2.0, a dataset containing biology protocols and corresponding pseudocodes, which supports LLMs in the formulation and evaluation of SPFT. Our work is extensible to the assessment of LLMs on SPFT across various domains and other fields that require protocol generation for specific goals.
@@ 6 Limitations
We recognize several limitations. The predefined actions may not encompass all actions performed in a biology labs. The definitions of predefined actions may be incomplete. To precisely define an action, it is necessary to define not only the function but also the function arguments. The number of protocols in Bioprot 2.0 may be insufficient for evaluation purposes. The performance of ProtoMed-LLM may decline outside of biology. Addressing this requires redefining domain-specific actions and exploring other LLMs for diverse fields. Future work should investigate these cross-disciplinary implications. LLMs are continuously evolving due to regular updates. The LLMs used for evaluation in this work might become unavailable in the future. Upgraded versions of LLMs may result in performance degradation and metrics may differ from those obtained using previous models. Due to selecting Llama-3 as the evaluator, our results may be susceptible to its biases and hallucinations. The outcomes when evaluated with models other than Llama-3 are unknown. Future work should investigate the outcomes using different LLMs as an evaluator. Using an API of LLMs as an evaluator, such GPT, is often not free of charge and can be costly. We used GPT-4 generated responses as a baseline; however, it may not accurately represent the ground truth. Future work should explore the implications of employing alternative resources (e.g., manually annotated pseudocodes, responses generated by other models) as the baseline. We observed basic actions classified as NoAction in minor cases. It has been reported that GPT prefers outputs from LLMs, which also produced our evaluation materials including all ground truth and target pseudocodes. This can potentially influence the scores. The four criteria mentioned in G-Eval may not sufficiently fulfill the role of evaluating protocols where real-world validation is crucial. Also, applying these criteria originally designed for summarization tasks may be inappropriate for evaluating SPFT. Even if the protocol pseudocode is successfully synthesized, real-world experiments may fail depending on the person performing the protocol or the condition of the physical equipment, especially in cases that are more complex than stem cell culture or require delicate manual work and experience.
@@ Acknowledgements
We would like to thank Karim Md.Adnan for assisting us with the action defining process. This research was supported by a KIST project (2E32351) and Bio-Cluster Industry Capacity Enhancement Project of Jeonbuk Technopark (JBTP)